OpenStudy (anonymous):

sec(pi/2 - x) = cscx true or false?

OpenStudy (anonymous):

True my sir

OpenStudy (jtvatsim):

It's helpful to realize this is just testing your knowledge of \[\cos(\pi/2 - x) = \sin(x)\] because sec and csc in the question can be rewritten as \[\frac{1}{\cos(\pi/2 - x)} = \frac{1}{\sin(x)}\]

OpenStudy (anonymous):

the problem is i dont even understand the basics :/ thank you though!! :D

OpenStudy (jtvatsim):

lol, well keep practicing. It will come! :)